OpenStudy (samigupta8):

if the equation x^2+2(a+1)x+9a-5=0 has only negative roots then a satisfies the relation a>=6 .. how?

hartnn (hartnn):

find 2 roots using quadratic formula, one is surely negative, which is other one ?

hartnn (hartnn):

-2a -2 -sqrt {...} is negative even if sqrt term becomes greater than -2a-2

OpenStudy (samigupta8):

\[x=\frac{ -2a-2\pm \sqrt{4a^2-24a+24}}{ ? 2}\]

hartnn (hartnn):

for -2a-2 + sqrt {...} term equate it to -2a-2 + sqrt {...} < 0

OpenStudy (samigupta8):

do u want me to aply mid term splitting in that square root term and get the value of a

OpenStudy (samigupta8):

i m getting 20<34 a

OpenStudy (raden):

the 3 rules bellow must be satisfied : (1) x1 + x2 < 0 (2) x1 * x2 > 0 (3) D >= 0 extra information (known) a >=0 with x1, x2 = roots D = discriminant

OpenStudy (raden):

oppss.. i mean a >=6

OpenStudy (samigupta8):

yaa bt how that's wht we have to prove

OpenStudy (raden):

just find solution for a which satisfies that equation. x^2+2(a+1)x+9a-5=0 (1) x1 + x2 < 0 -b/a < 0 -2(a+1) < 0 a + 1 > 0 a > -1

OpenStudy (raden):

can you solve for the 2nd case ? (2) x1 * x2 > 0

OpenStudy (samigupta8):

yaaa sure

OpenStudy (samigupta8):

a>5/9 and a>-1

OpenStudy (raden):

a > -1 in the 1st case yeah, the 2nd is right how the 3rd case ?

OpenStudy (samigupta8):

4a^2-24a+24>=0

OpenStudy (raden):

hmmm, i got different like yours. 4a^2 - 28a + 24>=0, it can simplied be a^2 - 7a + 6 >=0

OpenStudy (samigupta8):

yaa sorry urs is correct

OpenStudy (samigupta8):

now do we need to take intersection of al of them

OpenStudy (raden):

yes, exactly find the intersection of them

OpenStudy (samigupta8):

i got the answer thanks...

OpenStudy (raden):

very welcome, dear :)

OpenStudy (raden):

intersection : a > = 6, right ?